# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v1

@@ III.2.3 HHL and Psi-HHL results
We begin by conducting a study similar to the one shown in Figure 4 for the toy matrices. It is not practical to carry out an analysis similar to that in Figure 3, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon. For the purposes of our analysis in Figure 4, we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 5 repetitions. For this analysis, we consider Li2, RbH and CsH. The details are:
@@
    Li2: 21-qubit circuit with an $A$ matrix size of $64\times 64$ (built out of single and double excitations arising from 3 occupied and 8 virtual orbitals) in the STO-6G basis, with $\mathcal{\kappa}=64.07$.
@@
    RbH: 25-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 13 occupied and 3 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=153.17$.
  3. 3.
    CsH: 26-qubit calculations with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 18 occupied and 2 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=466.59$. Although RbH and CsH have the same $A$ matrix size, the latter involves a 26-qubit calculation because we set an additional ancilla for CsH due to its larger atomic number.
We note that for RbH, the same calculation would have incurred more qubits for VQE-UCCSD (32-qubit circuit) and QPE-CASCI (with $n_{r}=8$, 40-qubit circuit) computations, while we carry out a 25-qubit circuit evaluation in the case of HHL-LCC. Similarly, for CsH, VQE-UCCSD: 40 qubits, QPE-CASCI: 48 qubits, and HHL-LCC: 26 qubits.
Our results are presented in Figure 5, with the accompanying data given in Table A11 of the Appendix. For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots. The data indicates that for Li2, the Psi-HHL approach gives better average PFDs, especially in the regime of fewer number of shots (500 and 1000 shots). In the same regime, the spread in PFDs is also lesser for Psi-HHL. For the cases of RbH and CsH, as we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller.
@@ IV Conclusion
We propose the Psi-HHL algorithm to efficiently handle problems involving $A$ matrices with large condition numbers. The approach involves post-selecting 0 in HHL (the ‘wrong signal’), followed by post-selecting 1 from an HHL-like circuit that contains an additional mixing $2\times 2$ unitary (the ‘mixed signal’), and taking the difference between the wrong and the mixed signals to extract the correct signal.
We present arguments to show that the complexity of Psi-HHL is optimal in condition number for large $\mathcal{\kappa}$ regimes; the complexity of the algorithm scales as $\mathcal{\kappa}$. This is an improvement over earlier approaches in literature, such as HHL with amplitude amplification (scales quadratically, and incurs several additional gates due to its iterative nature) and HHL with variable time amplitude amplification (scales as $\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa}$). It is also important to note that while the optimal $\mathcal{\kappa}$ behaviour can be recovered using a discrete quantum adiabatic approach based quantum linear solver algorithm [12], we achieve the same result using Psi-HHL, which is a simple modification to the HHL algorithm.
We apply our method to different toy matrices, where we systematically vary the condition number all the way to about 1 million while keeping the number of shots fixed to 1 million, and show that while HHL fails beyond a certain value, Psi-HHL continues to predict good results. We also demonstrate the performance of our proposed Psi-HHL approach for quantum chemistry examples, where we once again find that Psi-HHL yields better quality results in much fewer number of shots. En route, we also carry out a qualitative analysis to check suitability of treating LCC equations using the HHL algorithm.
@@ Acknowledgements.
We acknowledge National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Siddhi-AI’, under National PARAM Supercomputing Facility (NPSF), C-DAC, Pune and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. VSP thanks Prof. Aram Harrow for a very fruitful discussion on HHL as well as HHL-LCC. VSP and AJ thank Dr. Shibdas Roy for nice discussions on complexity. VSP and PBT thank Prof. Debashis Mukherjee for useful discussions on the quantum chemical sections of the work. KS acknowledges support from Quantum Leap Fragship Program (Grant No. JPMXS0120319794) from Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan; Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from Japan Science and Technology Agency (JST), Japan; and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from Japan Society for the Promotion of Sciences (JSPS), Japan.
@@
Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy matrices for comparing the performance of HHL and the Psi-HHL algorithms. The cell with a star symbol denotes the matrix element with the variable, $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
@@
    $A$ is diagonal and the amplitude encoded state corresponding to $\vec{b}$ is in an equal superposition. In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is set to 1.
@@
Figure A2: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{5}$ shots throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Subfigures (g) through (i) consider the case where $A$ is not diagonal and $\vec{b}$ is in unequal superposition.
@@
Figure A3: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{6}$ throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal, and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition.
@@
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
[…]
$2^{2}$ 8 -0.44(0.65, 0.35) -0.22(0.21, 0.79) -1.01(0.92, 0.08) -0.10(0.06, 0.94) -0.87(0.92, 0.08) -0.04(0.08, 0.92)
$2^{3}$ 9 -0.55(0.72, 0.28) -0.07(0.27, 0.73) -5.82(0.97, 0.03) -0.32(0.15, 0.85) -5.72(0.97, 0.03) 0.08(0.15, 0.85)
$2^{4}$ 10 -0.81(0.74, 0.26) 0.32(0.32, 0.68) 14.39(0.99, 0.01) -0.69(0.2, 0.80) 17.20(0.99, 0.01) 0.26(0.2, 0.80)
$2^{5}$ 11 -0.17(0.75, 0.25) -0.02(0.35, 0.65) 3.96(0.99, 0.01) 2.85(0.23, 0.77) -0.12(0.99, 0.01) 0.32(0.22, 0.78)
$2^{6}$ 12 0.30(0.75, 0.25) 0.05(0.37, 0.63) -6.73(0.99, 0.01) 1.66(0.24, 0.76) -6.73(0.99, 0.01) -1.56(0.24, 0.76)
$2^{7}$ 13 nan(0.75, 0.25) -0.45(0.37, 0.63) nan(0.99, 0.01) 1.91(0.25, 0.75) nan(0.99, 0.01) -2.62(0.25, 0.75)
$2^{8}$ 14 nan(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) -2.00(0.25, 0.75) nan(0.99, 0.01) -3.87(0.25, 0.75)
$2^{9}$ 15 nan(0.75, 0.25) -0.22(0.37, 0.63) nan(0.99, 0.01) -7.50(0.25, 0.75) nan(0.99, 0.01) 8.02(0.25, 0.75)
$2^{10}$ 16 nan(0.75, 0.25) -0.26(0.37, 0.63) nan(0.99, 0.01) -2.33(0.25, 0.75) nan(0.99, 0.01) -7.77(0.25, 0.75)
$2^{11}$ 17 nan(0.75, 0.25) 0.08(0.37, 0.63) nan(0.99, 0.01) 2.65(0.25, 0.75) nan(0.99, 0.01) -0.55(0.25, 0.75)
$2^{12}$ 18 nan(0.75, 0.25) 0.03(0.37, 0.63) nan(0.99, 0.01) 7.93(0.25, 0.75) nan(0.99, 0.01) -2.41(0.25, 0.75)
$2^{13}$ 19 nan(0.75, 0.25) 0.30(0.38, 0.62) nan(0.99, 0.01) 4.67(0.25, 0.75) nan(0.99, 0.01) 3.12(0.25, 0.75)
$2^{14}$ 20 nan(0.75, 0.25) 0.41(0.38, 0.62) nan(0.99, 0.01) 9.11(0.25, 0.75) nan(0.99, 0.01) -5.60(0.25, 0.75)
$2^{15}$ 21 nan(0.75, 0.25) 0.13(0.38, 0.62) nan(0.99, 0.01) 9.42(0.26,0.74) nan(0.99, 0.01) -1.11(0.25, 0.75)
$2^{16}$ 22 nan(0.75, 0.25) 0.54(0.38, 0.62) nan(0.99, 0.01) -0.25(0.25, 0.75) nan(0.99, 0.01) 2.43(0.25, 0.75)
$2^{17}$ 23 nan(0.75, 0.25) -0.03(0.38, 0.62) nan(0.99, 0.01) -0.40(0.26, 0.74) nan(0.99, 0.01) -7.87(0.25, 0.75)
$2^{18}$ 24 nan(0.75, 0.25) 0.26(0.37, 0.63) nan(0.99, 0.01) -8.56(0.26, 0.74) nan(0.99, 0.01) 9.69(0.26, 0.74)
$2^{19}$ 25 nan(0.75, 0.25) 0.19(0.37, 0.63) nan(0.99, 0.01) 6.91(0.26, 0.74) nan(0.99, 0.01) 1.06(0.25, 0.75)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.58(0.37, 0.63) nan(0.99, 0.01) 0.30(0.25, 0.75) nan(0.99, 0.01) 6.15(0.25, 0.75)
$2^{2}$ 8 -0.57(0.65, 0.35) 0.00(0.21, 0.79) -0.87(0.92, 0.08) 0.03(0.08, 0.92) -0.88(0.92, 0.08) 0.04(0.08, 0.92)
$2^{3}$ 9 -0.61(0.72, 0.28) -0.14(0.27, 0.73) -5.82(0.97, 0.03) -0.04(0.15, 0.85) -5.82(0.97, 0.03) 0.00(0.15, 0.85)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.04(0.31, 0.69) 14.66(0.99, 0.01) 0.11(0.20, 0.8) 14.95(0.99, 0.01) -0.15(0.20, 0.80)
$2^{5}$ 11 0.27(0.75, 0.25) -0.06(0.34, 0.66) 1.84(0.99, 0.01) 0.16(0.22, 0.78) 0.61(0.99, 0.01) 0.44(0.22, 0.78)
$2^{6}$ 12 0.04(0.75, 0.25) 0.00(0.36, 0.64) -0.72(0.99, 0.01) 0.01(0.24, 0.76) 1.68(0.99, 0.01) -0.70(0.24, 0.76)
$2^{7}$ 13 -0.07(0.75, 0.25) 0.20(0.37, 0.63) 1.35(0.99, 0.01) -1.56(0.25, 0.75) -4.87(0.99, 0.01) -0.11(0.25, 0.75)
$2^{8}$ 14 -0.13(0.75, 0.25) -0.02(0.37, 0.63) 0.09(0.99, 0.01) -0.04(0.25, 0.75) -4.04(0.99, 0.01) -1.15(0.25, 0.75)
$2^{9}$ 15 0.05(0.75, 0.25) 0.06(0.37, 0.63) -2.82(0.99, 0.01) 2.82(0.25, 0.75) nan(0.99, 0.01) 2.26(0.25, 0.75)
$2^{10}$ 16 0.04(0.75, 0.25) -0.04(0.37, 0.63) nan(0.99, 0.01) -0.12(0.25, 0.75) nan(0.99, 0.01) -0.59(0.25, 0.75)
$2^{11}$ 17 0.19(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -1.49(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
$2^{12}$ 18 0.08(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) 0.11(0.25, 0.75) nan(0.99, 0.01) 1.26(0.25, 0.75)
$2^{13}$ 19 0.12(0.75, 0.25) 0.01(0.37, 0.63) nan(0.99, 0.01) 0.13(0.25, 0.75) nan(0.99, 0.01) -3.37(0.25, 0.75)
$2^{14}$ 20 -0.03(0.75, 0.25) -0.00(0.37, 0.63) nan(0.99, 0.01) -4.52(0.25, 0.75) nan(0.99, 0.01) 1.47(0.25, 0.75)
$2^{15}$ 21 0.05(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -0.70(0.25, 0.75) nan(0.99, 0.01) -0.79(0.25, 0.75)
$2^{16}$ 22 0.10(0.75, 0.25) 0.12(0.38, 0.63) nan(0.99, 0.01) 1.57(0.25, 0.75) nan(0.99, 0.01) 0.84(0.25, 0.75)
$2^{17}$ 23 -0.03(0.75, 0.25) -0.15(0.38, 0.63) nan(0.99, 0.01) -0.40(0.25, 0.75) nan(0.99, 0.01) 0.05(0.25, 0.75)
$2^{18}$ 24 0.11(0.25, 0.75) 0.00(0.37, 0.63) nan(0.99, 0.01) -1.22(0.25, 0.75) nan(0.99, 0.01) -3.58(0.25, 0.75)
$2^{19}$ 25 0.08(0.75, 0.25) -0.03(0.37, 0.63) nan(0.99, 0.01) 0.78(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.19(0.37, 0.63) nan(0.99, 0.01) -1.52(0.25, 0.75) nan(0.99, 0.01) 0.01(0.25, 0.75)

### v2

@@ III.2.3 HHL and Psi-HHL results
We begin by conducting a study similar to the one shown in Figure 5 for the toy matrices. It is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon. We next move to Figure 7 (with the accompanying data given in Table A12 of the Appendix), where we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions. For this analysis, we consider Li2 and KH. The details are:
@@
    Li2: 21-qubit circuit with an $A$ matrix size of $64\times 64$ (built out of single and double excitations arising from 3 occupied and 7 virtual orbitals) in the STO-6G basis, with $\mathcal{\kappa}=64.07$.
@@
    KH: 26-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 6 occupied and 6 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=365.40$.
We pick these two systems, as in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300. Repeating the calculations for RbH and CsH may not give any new insights, and hence we restrict the results only to Li2 and KH. For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots. The data clearly shows that for Li2, the Psi-HHL approach always has a smaller spread in PFDs when compared to HHL. For KH, we observe that at 1000 shots, HHL fails, while Psi-HHL does not, but comes with large error bars. As we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller, but Psi-HHL consistently has smaller error bars than HHL.
We now move to our next set of results, namely the PEC for the KH molecule. Since it is computationally not feasible to pick the larger system sizes from Figure 6, we choose the 6 occupied and 6 virtual orbitals case. This choice corresponds to a 26-qubit calculation with an $A$ matrix size of $256\times 256$. We consider geometries in the neighbourhood of the equilibrium bond length (3 geometries in steps of 0.47 Bohr on either side of the equilibrium bond length). With such a choice, the correlation energies required to be captured ranges from about 0.3 mHa (milli-Hartree) for 5.65 Bohr geometry to 3 mHa for the 2.82 Bohr geometry (the two extreme points in the 7 geometries considered), thus allowing us to check the performance of HHL and Psi-HHL across a range of correlation energy values. Calculating energies across a range of bond lengths further allows us to check across a (limited) range of condition numbers (the smallest being 365.40 and the largest 393.21) the behaviour of HHL and Psi-HHL. Each calculation involves 50000 shots and our reported results are averaged over 30 repetitions. For our Psi-HHL computations, we pick $\alpha$ of 60 degrees. Figure 8(a) (and the accompanying Table A13 of the Appendix) shows the results for $\Delta E$, which is the difference between HHL (Psi-HHL) total energy and and the classical LCC value for the considered geometries. We find that Psi-HHL performs better than HHL in predicting correlation energies, especially on the extreme ends of our PEC when $\mathcal{\kappa}$ values are relatively larger than for those points around the equilibrium geometry. When we examine Figure 8(b), we see that Psi-HHL performs consistently better than HHL. Furthermore, the spread in results across 30 repetitions follows a trend globally: it increases from 2.82 through 5.65 Bohr calculations. We recall that the correlation energies themselves decrease monotonically and by as much as an order of magnitude as we move from 2.82 Bohr to 5.65 Bohr geometries. Thus, although the condition numbers are comparable (for example, 374.93 and 371.38 for the 3.29 and the 4.71 Bohr cases respectively), since the quantity to be captured (the correlation energy) itself is much smaller for the latter, we need to supply more shots to restrict the spread in PFD across repeats. However, since we keep the number of shots fixed throughout the PEC, we observe a larger spread in PFDs. We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using HHL, with 1000 shots and 20 repeats. We average our results over 20 repetitions. $a$ is varied from 0.26 to 0.99 (below 0.26, we get nan). We observe the same trend as in Figure 8(b) (see Figure A10 of Appendix).
@@ IV Conclusion
We propose the Psi-HHL algorithm to efficiently handle problems involving $A$ matrices with large condition numbers ($\mathcal{\kappa}$). The approach involves post-selecting 0 in HHL (the ‘wrong signal’), followed by post-selecting 1 from an HHL-like circuit that contains an additional mixing $2\times 2$ unitary (the ‘mixed signal’), and taking the difference between the wrong and the mixed signals to extract the correct signal.
We present arguments to show that the complexity of Psi-HHL is optimal in condition number for large $\mathcal{\kappa}$ regimes; the complexity of the algorithm scales as $\mathcal{O}(\mathcal{\kappa})$. This is an improvement over earlier approaches in literature, such as HHL with amplitude amplification (scales quadratically, and incurs several additional gates due to its iterative nature) and HHL with variable time amplitude amplification (scales as $\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa}$). It is also important to note that while the optimal $\mathcal{\kappa}$ behaviour can be recovered using a discrete quantum adiabatic approach based quantum linear solver algorithm [12], we achieve the same result using Psi-HHL, which is a simple modification to the HHL algorithm. We also discuss the case of applying Psi-HHL to singular matrices.
We apply our method to different toy matrices, where we systematically vary the condition number all the way to about 1 million and keeping the number of shots fixed to 1 million for each of the computations, and show that while HHL fails beyond a certain value of $\mathcal{\kappa}$, Psi-HHL continues to predict good results that we quantify via percentage fraction difference (PFD) relative to the classical value as well as via spread in the PFD across many repetitions. We also demonstrate the performance of our proposed Psi-HHL approach for quantum chemistry examples with up to 26-qubit computations, and we once again find that Psi-HHL yields better quality results in much fewer number of shots. En route, we also carry out a qualitative analysis to check suitability of treating LCC equations using the HHL algorithm.
@@ Acknowledgements.
We acknowledge National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Siddhi-AI’, under National PARAM Supercomputing Facility (NPSF), C-DAC, Pune and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. VSP thanks Prof. Aram Harrow for a very fruitful discussion on HHL as well as HHL-LCC. VSP and AJ thank Dr. Shibdas Roy for nice discussions on complexity. PBT thank Prof. Debashis Mukherjee for useful discussions on the quantum chemical sections of the work. KS acknowledges support from Quantum Leap Fragship Program (Grant No. JPMXS0120319794) from Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan; Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from Japan Science and Technology Agency (JST), Japan; and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from Japan Society for the Promotion of Sciences (JSPS), Japan.
@@
Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy $A$ matrices and un-normalized $\vec{b}$ for comparing the performance of HHL and the Psi-HHL algorithms. The cell with a star symbol denotes the matrix element with the variable $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$. A white cell denotes a zero matrix element.
@@
    $A$ is diagonal and the amplitude encoded state corresponding to $\vec{b}$ is in an equal superposition. In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, and the un-normalized $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is set to 1.
@@
Figure A2: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{5}$ shots throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal and $\vec{b}$ is in unequal superposition. Each data point is generated by taking the average over 50 repetitions.
@@
Figure A3: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{6}$ throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal, and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition. Each data point is generated by taking the average over 50 repetitions.
@@
$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
# shots $\approx\kappa$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.08(0.64, 0.36) 0.01(0.21, 0.79) 0.03(0.92, 0.08) -0.11(0.08, 0.92) 0.07(0.92, 0.08) 0.01(0.08, 0.92)
$2^{3}$ 9 0.01(0.72, 0.28) -0.17(0.27, 0.73) -0.05(0.97, 0.03) -0.09(0.15, 0.85) -0.02(0.97, 0.03) -0.05(0.15, 0.85)
$2^{4}$ 10 -0.01(0.74, 0.26) 0.06(0.32, 0.68) 0.31(0.99, 0.01) 0.47(0.20, 0.80) 0.48(0.99, 0.01) 0.39(0.20, 0.80)
$2^{5}$ 11 -0.06(0.75, 0.25) -0.14(0.35, 0.65) 1.85(0.99, 0.01) -0.04(0.23, 0.77) -0.10(0.99, 0.01) -0.15(0.22, 0.78)
$2^{6}$ 12 0.08(0.75, 0.25) -0.38(0.37, 0.63) nan(0.99, 0.01) -1.41(0.24, 0.76) nan(0.99, 0.01) -1.75(0.24, 0.76)
$2^{7}$ 13 0.00(0.75, 0.25) 0.10(0.37, 0.63) nan(0.99, 0.01) 1.39(0.25, 0.75) nan(0.99, 0.01) 1.85(0.25, 0.75)
$2^{8}$ 14 0.14(0.75, 0.25) -0.10(0.37, 0.63) nan(0.99, 0.01) 0.42(0.25, 0.75) nan(0.99, 0.01) 1.92(0.25, 0.75)
$2^{9}$ 15 0.12(0.75, 0.25) 0.14(0.37, 0.63) nan(0.99, 0.01) 3.21(0.25, 0.75) nan(0.99, 0.01) -1.05(0.25, 0.75)
$2^{10}$ 16 -0.01(0.75, 0.25) -0.08(0.37, 0.63) nan(0.99, 0.01) 0.87(0.25, 0.75) nan(0.99, 0.01) 1.00(0.25, 0.75)
$2^{11}$ 17 -0.05(0.75, 0.25) -0.28(0.37, 0.63) nan(0.99, 0.01) -0.23(0.25, 0.75) nan(0.99, 0.01) 2.19(0.25, 0.75)
$2^{12}$ 18 0.02(0.75, 0.25) -0.03(0.37, 0.63) nan(0.99, 0.01) 1.38(0.25, 0.75) nan(0.99, 0.01) 0.96(0.25, 0.75)
$2^{13}$ 19 0.09(0.75, 0.25) -0.38(0.38, 0.62) nan(0.99, 0.01) -0.59(0.25, 0.75) nan(0.99, 0.01) -2.71(0.25, 0.75)
$2^{14}$ 20 0.27(0.75, 0.25) 0.03(0.38, 0.62) nan(0.99, 0.01) 2.97(0.25, 0.75) nan(0.99, 0.01) -0.36(0.25, 0.75)
$2^{15}$ 21 0.21(0.75, 0.25) -0.09(0.38, 0.62) nan(0.99, 0.01) -3.51(0.26,0.74) nan(0.99, 0.01) 2.68(0.25, 0.75)
$2^{16}$ 22 -0.11(0.75, 0.25) 0.05(0.38, 0.62) nan(0.99, 0.01) -1.74(0.25, 0.75) nan(0.99, 0.01) 1.15(0.25, 0.75)
$10^{5}$ $2^{17}$ 23 0.29(0.75,0.25) 0.05(0.38, 0.62) nan(0.99, 0.01) 3.06(0.26, 0.74) nan(0.99, 0.01) 2.17(0.25, 0.75)
$2^{2}$ 8 0.00(0.64, 0.36) 0.05(0.21, 0.79) 0.03(0.92, 0.08) 0.02(0.08, 0.92) 0.05(0.92, 0.08) -0.03(0.08, 0.92)
$2^{3}$ 9 0.00(0.72, 0.28) -0.05(0.27, 0.73) 0.02(0.97, 0.03) 0.00(0.15, 0.85) -0.08(0.97, 0.03) 0.01(0.15, 0.85)
$2^{4}$ 10 0.00(0.74, 0.26) 0.04(0.27, 0.73) -0.07(0.99, 0.01) -0.10(0.20, 0.8) 0.12(0.99, 0.01) 0.14(0.20, 0.80)
$2^{5}$ 11 0.06(0.75, 0.25) -0.02(0.31, 0.69) 0.12(0.99, 0.01) 0.07(0.22, 0.78) 0.16(0.99, 0.01) 0.37(0.22, 0.78)
$2^{6}$ 12 0.06(0.75, 0.25) 0.05(0.34, 0.66) 0.99(0.99, 0.01) 0.29(0.24, 0.76) 0.02(0.99, 0.01) -0.03(0.24, 0.76)
$2^{7}$ 13 0.10(0.75, 0.25) -0.03(0.36, 0.64) 1.36(0.99, 0.01) 0.17(0.25, 0.75) 2.19(0.99, 0.01) 0.50(0.25, 0.75)
$2^{8}$ 14 -0.07(0.75, 0.25) 0.07(0.37, 0.63) 8.58(0.99, 0.01) -0.03(0.25, 0.75) 3.63(0.99, 0.01) -1.34(0.25, 0.75)
$2^{9}$ 15 0.04(0.75, 0.25) 0.04(0.37, 0.63) nan(0.99, 0.01) -0.05(0.25, 0.75) nan(0.99, 0.01) 0.25(0.25, 0.75)
$2^{10}$ 16 -0.14(0.75, 0.25) 0.05(0.37, 0.63) nan(0.99, 0.01) -0.25(0.25, 0.75) nan(0.99, 0.01) 0.71(0.25, 0.75)
$2^{11}$ 17 -0.09(0.75, 0.25) 0.10(0.37, 0.63) nan(0.99, 0.01) -0.03(0.25, 0.75) nan(0.99, 0.01) -0.31(0.25, 0.75)
$2^{12}$ 18 -0.03(0.75, 0.25) 0.06(0.37, 0.63) nan(0.99, 0.01) 0.78(0.25, 0.75) nan(0.99, 0.01) -0.82(0.25, 0.75)
$2^{13}$ 19 0.04(0.75, 0.25) -0.02(0.37, 0.63) nan(0.99, 0.01) 0.75(0.25, 0.75) nan(0.99, 0.01) -0.87(0.25, 0.75)
$2^{14}$ 20 -0.12(0.75, 0.25) 0.09(0.37, 0.63) nan(0.99, 0.01) 0.07(0.25, 0.75) nan(0.99, 0.01) -0.05(0.25, 0.75)
$2^{15}$ 21 0.02(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) 1.00(0.25, 0.75) nan(0.99, 0.01) 0.14(0.25, 0.75)
$2^{16}$ 22 0.09(0.75, 0.25) 0.02(0.37, 0.63) nan(0.99, 0.01) -0.09(0.25, 0.75) nan(0.99, 0.01) 0.21(0.25, 0.75)
$10^{6}$ $2^{17}$ 23 0.02(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) -0.66(0.25, 0.75) nan(0.99, 0.01) 0.45(0.25, 0.75)